Find the general/particular solution of the following Differential Equations.


(Integrable Combinations)


y(y² + 1)dx + x(y² - 1)dy=0


1
Expert's answer
2021-09-14T00:00:20-0400

Solution

Let’s rewrite eq. in the form

"y^3dx+xy^2dy+ydx-xdy=0"

Dividing by y2 we obtain

"ydx+xdy+\frac{ydx-xdy}{y^2}=0"

According to Product Rule and Quotient Rule

"d\left(xy\right)+d\left(\frac{x}{y}\right)=0"

Thus integrating

"xy+\frac{x}{y}=C"

Here C is arbitrary constant.

Answer: xy+x/y=C